# A Scalable Quantum Non-local Neural Network for Image Classification

### v1

@@ Conclusion
Loss Convergence Analysis
In the QNL-Net framework, a hybrid gradient backpropagation approach is used to train our model effectively. This approach comprises optimizing both the classical parameters in neural nets and the set of trainable parameters, which are the angles of quantum gates in VQCs. This hybrid training approach first applies a forward pass to optimize parameters and converge the loss function. Our model uses the negative log-likelihood (NLL) loss function for the binary classification problem. The NLL loss measures the variation between the true labels $y$ and the classical predicted probabilities $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]=p(y|x;\theta)$ obtained from the measurement of the hybrid classical-QNL-Net model, and is defined for binary classification as:
$\mathcal{L}(\theta,\phi)=-\sum^{n}_{i=1}\left(y_{i}\log\hat{y}_{i1}+(1-y_{i})\log\hat{y}_{i0}\right),$ (40)
where $n$ is the number of samples in the dataset, $\hat{y}_{0}$ is the predicted probability for class 0, $\hat{y}_{1}$ is the predicted probability for class 1, $\theta$ are the quantum parameters, and $\phi$ are the classical parameters.
To optimize both classical and quantum parameters, we compute the gradients of the loss with respect to the predicted probabilities,
@@
Then, we compute the gradient of the predicted probabilities with respect to the quantum parameters using the parameter-shift rule (Mitarai et al. 2018; Schuld et al. 2019) to obtain,
$\frac{\partial\hat{y}_{i1}}{\partial\theta_{j}}=\frac{p(y_{i}|x;\theta_{j}+\frac{\pi}{2})-p(y_{i}|x;\theta_{j}-\frac{\pi}{2})}{2}.$ (42)
Now, we can derive the gradient of the loss function with respect to the quantum parameters from eq.(41) & (42), such that,
$\frac{\partial\mathcal{L}}{\partial\theta_{j}}=\sum_{i}\left(\frac{\partial\mathcal{L}}{\partial\hat{y}_{i1}}\frac{\partial\hat{y}_{i1}}{\partial\theta_{j}}+\frac{\partial\mathcal{L}}{\partial\hat{y}_{i0}}\frac{\partial\hat{y}_{i0}}{\partial\theta_{j}}\right).$ (43)
For the classical parameters, we calculate the gradients of the predicted probabilities using standard backpropagation techniques,
@@
where $h$ denotes the output from the classical model. These derived first-order objective functions are optimized using the stochastic gradient descent method, Adam (Kingma and Ba 2014). The first-order moment $m_{t}$ and the second-order moment for the quantum gradient (from eq.(43)) are estimated using the following equations,
$m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\frac{\partial\mathcal{L}}{\partial\theta_{j}},$ (45)
$v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\left(\frac{\partial\mathcal{L}}{\partial\theta_{j}}\right)^{2},$ (46)
where $t$ is the iteration/time-step and constants $\beta_{1}$ $\beta_{2}$ are the exponential decay rate. These moments are corrected for initialization bias and we obtain bias-corrected moments, such that,
@@
$\theta_{j}\leftarrow\theta_{j}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon},$ (48)
where $\eta$ is the learning rate and $\epsilon$ is an added small constant for numerical stability. The optimization process is similar to the above for classical parameters $\phi$ for the classical gradient in eq.(44).
@@
Datasets
MNIST (Deng 2012) is a handwritten digit recognition dataset used for many machine learning and computer vision tasks. Each image in MNIST is a grayscale 28 x 28-pixel representation of handwritten digits ranging from 0 to 9. The MNIST dataset contains 60,000 training samples used to train models and 10,000 testing samples used to evaluate model performance. These samples are handwritten by various individuals, covering a lot of variations and styles, ideal for machine learning.
CIFAR-10 (Krizhevsky, Hinton et al. 2009) is another widely-used benchmark dataset in the field of computer vision. It presents a collection of 32 x 32 size RGB images distributed across ten classes, including images of objects such as airplanes, cars, birds, cats, etc. The dataset contains a total of 50,000 training samples (5000 training samples per class) and 10,000 testing samples (1000 testing samples per class). Its diverse set of classes, coupled with variations in lighting, angle, and pose within images, makes it a suitable dataset for evaluating the robustness and generalization capability of image classification models.
Figure 5: Sample images from the datasets used in this work: (a) MNIST (Deng 2012) classes 0 and 1; and (b) CIFAR-10 (Krizhevsky, Hinton et al. 2009) classes 2 and 8, i.e., bird and ship respectively.

### v2

@@ Conclusion
Gradient Calculations
In the proposed QNL-Net framework, a hybrid gradient backpropagation approach is used to train our model effectively. This approach comprises optimizing both the classical parameters present in the neural nets and the set of trainable classical parameters which are the angles of quantum gates in the VQC. This hybrid training approach first applies a forward pass to optimize parameters for the convergence of the loss function. Our model uses the negative log-likelihood (NLL) loss function for the binary classification problem. The NLL loss measures the variation between the true labels $y$ and the classical predicted probabilities $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]$ obtained from the measurement of the hybrid quantum-classical QNL-Net model, and is defined for binary classification as:
$\mathcal{L}(\phi)=-\sum^{N}_{i=1}\left(y_{i}\log\hat{y}_{i1}+(1-y_{i})\log\hat{y}_{i0}\right),$ (43)
where $N$ is the number of samples in the dataset, $\hat{y}_{0}$ is the predicted probability for class 0, $\hat{y}_{1}$ is the predicted probability for class 1, and $\phi$ are the classical parameters.
To optimize these parameters, we compute the gradients of the loss with respect to the predicted probabilities,
@@
The gradients of the predicted probabilities are then computed using standard back-propagation techniques as follows,
@@
where $h$ denotes the output from the classical model.
These derived first-order objective functions are optimized using the stochastic gradient descent method, Adam (Kingma and Ba 2014). The first-order moment $m_{t}$ and the second-order moment $v_{t}$ for the gradient (from Eq. (45)) are estimated using the following equations,
$m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\frac{\partial\mathcal{L}}{\partial\phi},$ (46)
$v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\left(\frac{\partial\mathcal{L}}{\partial\phi}\right)^{2},$ (47)
where $t$ is the iteration/time-step and constants $\beta_{1}$ $\beta_{2}$ are the exponential decay rate. These moments are corrected for initialization bias, and we obtain bias-corrected moments such that,
@@
$\phi\leftarrow\phi-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon},$ (49)
where $\eta$ is the learning rate and $\epsilon$ is an added small constant for numerical stability.
@@
Datasets Utilized
MNIST (Lecun et al. 1998) is a handwritten digit recognition dataset used for many machine learning and computer vision tasks. Each image in MNIST is a grayscale 28 x 28-pixel representation of handwritten digits ranging from 0 to 9. The MNIST dataset contains 60,000 training samples used to train models and 10,000 testing samples used to evaluate model performance. These samples are handwritten by various individuals, covering a lot of variations and styles, ideal for machine learning.
CIFAR-10 (Krizhevsky, Sutskever, and Hinton 2012) is another widely-used benchmark dataset in the field of computer vision. It presents a collection of 32 x 32 size RGB images distributed across ten classes, including images of objects such as airplanes, cars, birds, cats, etc. The dataset contains a total of 50,000 training samples (5000 training samples per class) and 10,000 testing samples (1000 testing samples per class). Its diverse set of classes, coupled with variations in lighting, angle, and pose within images, makes it a suitable dataset for evaluating the robustness and generalization capability of image classification models.
Figure 5: Sample images from the datasets used: MNIST (Lecun et al. 1998) (classes 0 and 1) and CIFAR-10 (Krizhevsky, Sutskever, and Hinton 2012) (classes 2 and 8, i.e., bird and ship respectively).